# MergeMix: A Unified Augmentation Paradigm for Visual and Multi-Modal Understanding

### v2

@@ 5.3 Implementations
In this subsection, we briefly introduce the implementations on the classification task and the MLLM benchmark. The full description is in Appendix B.2. For the CIFAR100, images are resized to $224\times 224$ for ViT-based models (e.g., DeiT) and trained with AdamW (weight decay 0.05), batch size 100, for 200 or 600 epochs. We use RandomFlip and RandomCrop, plus RandAugment (cubuk2020randaugment). Learning rates are 1e-3 (DeiT-Tiny/Small, cosine schedule), 5e-4 (ViT-Small/Base), and 2e-4 (ViT-Large). For ImageNet-1K, we adopt the same settings but use a 1e-3 learning rate, batch size 1024, and 300 epochs for DeiT-Tiny/Small. For fine-grained datasets (CUB200, FGVC-Aircrafts, Stanford-Cars), we fine-tune DeiT-Small and ViT-Base for 200 epochs, batch size 16, learning rate 1e-5, using PyTorch pretrained weights (nips2019pytorch).
Following LLaVA-v1.5, we adopt Vicuna-v1.5 7B (chiang2023vicuna) as the language decoder and use a pre-trained 2-layer MLP projection to align visual and textual modalities, trained for one epoch on LCS-558K. The vision encoder is a pre-trained CLIP model that extracts image representations. During SFT, we train for one epoch on llava-v1.5-mix665k with a 2e-5 learning rate, batch size 64, and AdamW optimizer, using a 0.03 warmup ratio and cosine scheduler. Unlike LLaVA, we unfreeze the vision encoder during training. For Qwen2.5-VL-Instruction, we post training within the official checkpoints and the llava-v1.5-mix665k dataset for 0.1 epoch under similar optimization settings. The learning rates for the vision encoder, LLM decoder, and merger are set to 2e-6, 1e-5, and 1e-5, respectively.
@@ 5.6 Results of Calibration
DNNs are prone to overconfidence in classification tasks. Mainfold Mixup (verma2019manifold) found that the mixup methods can effectively alleviate this problem. To this end, we compute the Expected Calibration Error (ECE (guo2017calibration)) of various mixup approaches on the CIFAR100 dataset for image classification. Also, to further analyze the calibration of MLLMs, we implement four short answer tasks, POPE, GQA, ScienceVQA, and SEED. Figure 4 shows the results of DeiT-Tiny and ViT-Small models trained for 200 epochs, showing that MergeMix obtains the best calibration of 6.7% and 9.7% in those ViT-specific mixup methods. (i.e., TransMix, SMMix, and MixPro). Table 6 shows the results of the LLaVA baseline, LLaVA with SFT, and LLaVA with our approach. SFT reduces the ECE when tuning the vision encoder, with augmentation and ranking loss, which can be better since we bring in the reward signal for the model. The more comprehensive results of CIFAR100 and MLLM benchmark we plot in Table A3 and Table A9.
@@ B.1 Datasets
In this subsection, we will describe the datasets we chose in detail.
@@
We choose five mainstream classification datasets: (i) CIFAR100 dataset (krizhevsky2009learning) consists of 100 classes of color images with a resolution of 32 $\times$ 32 pixels, containing 50,000 training images and 10,000 test images. (ii) ImageNet-1k dataset (russakovsky2015imagenet) consists of 1,000 classes with varied image resolutions commonly cropped to 224 $\times$ 224 pixels, containing 1,281,167 training images and 50,000 test images. (iii) CUB200 dataset (wah2011caltech) contains 200 bird species, including total 11,788 images, we divided 5,994 images as training images and 5,794 images as test images, (iv) FGVC-Aircrafts dataset (maji2013fine) contains 100 aircraft model classes, including 6,667 training images and 3,333 test images, (v) Stanford-Cars dataset (jonathan2013cars) contains 196 car model classes, including 8,144 training images and 8,041 test images. All the fine-grained datasets we used set the resolution as 224 $\times$ 224 pixels for training and testing.
@@ B.2 Implementations
(i) For the CIFAR100 dataset, aiming to be suitable for training ViT-based approaches, e.g., DeiT, we resize images to $224\times 224$ and train them with the AdamW optimizer with weight decay of 0.05, batch size of 100, and total training of 200 epochs and 600 epochs. Uses RandomFlip and RandomCrop as basic augmentations, and additionally, we use RandAugment (cubuk2020randaugment). For DeiT-Tiny and DeiT-Small, we use the learning rate of 1e-3 with a dynamic cosine scheduler. For ViT-Small and ViT-Base models, we set the learning rate to 5e-4, the learning rate of ViT-Large up to 2e-4, all dynamically adjusted by a cosine scheduler. (ii) For the ImageNet-1K dataset, the dataset settings are the same as CIFAR100, but we use the learning rate as 1e-3, batch size of 1024, and a total training of 300 epochs for DeiT-Tiny and DeiT-Small with AdamW optimizer with weight decay of 0.05. (iii) For all fine-grain datasets, i.e., CUB-200 dataset, FGVC-Aircrafts dataset, and Stanford-Cars dataset, we fine-tune the DeiT-Small and ViT-Base model for 200 epochs with a batch size of 16, learning rate of 1e-5, loading the pre-trained model weight from PyTorch (nips2019pytorch).
Table A1: Top-1 accuracy (%) of mixup methods on CIFAR-100 dataset under DeiT-Tiny/Small, ViT-Small/Base/Large different model sizes. The $\alpha$ parameter of the Beta distribution follows the setting in OpenMixup (2022openmixup) setting.
@@
Following the LLaVA-v1.5 settings, we use a pre-trained Vicuna-v1.5 7B (chiang2023vicuna) as the language decoder, which uses a pre-trained 2 $\times$ MLP as the projection for aligning the vision and text modistes, which was trained for one epoch on LCS-558K. For the vision encoder, we use a pre-trained CLIP encoder and extract the visual representation from the input images. For SFT, the learning rate was set as 2e-5, the batch size was 64, and training one epoch on llava-v1.5-mix665k dataset, uses AdamW optimizer with (0.9, 0.999) betas and epsilon of 1e-8, warmup ratio of 0.03 with a cosine scheduler. The difference from LLaVA is that we unfreeze the vision encoder during training. About Qwen2.5-VL-Instruction, we fine-tune with llava-v1.5-mix665k dataset for 0.1 epoch, uses AdamW optimizer with (0.9, 0.999) betas and epsilon of 1e-8 like LLaVA, warmup with 0.03 ratio. The learning rate of the vision encoder, LLM decoder, and merger were set to 2e-6, 1e-5, and 1e-5, respectively.
@@ B.4 Future work
There are still some shortcomings in MergeMix for MLLMs. In future work, we will explore this from two levels: (1) From the data perspective: MergeMix focuses on the enhancement of the image modality during training, while text inputs still remain raw. How to extend mixup to the text modality in MLLM tasks needs to be solved, as this can provide more fine-grained optimization guidelines. (2) From the model perspective: The token merging is still static and unlearned. Improving the merging strategy to make it learnable via metrics or backpropagation could enhance the token merging ability for mixing.
@@ C.2 Robustness Experiments of Mixup Augmentations
For further exploring the effectiveness of MergeMix, we applied our approach on a hyper-model MetaFormer (CAFormer) (yu2022metaformer) and two audio datasets for classification by HuBERT-Base (hsu2021hubert). ESC-50 (piczak2015esc) consisted of 50 classes, with 1,200 training samples and 400 validation samples, and a maximum duration of 3 seconds. UrbanSound8k (salamon2014urban8k) is a classification dataset consisting of 10 classes, with a maximum duration of 4 seconds, containing 7,079 training samples and 816 validation samples. Following the USB experimental setup, we fine-tuned the model for 100 epochs using the AdamW optimizer. The base learning rate was set to 1e-4, texttt5e-4, with a batch size of 32 and a weight decay coefficient of 5e-4. Both the compared shuffling methods and our proposed MergeMix can be directly transferred to audio data (treated as one-dimensional sequences). Table A8 below shows our reproduced comparison results on two audio datasets. Compared to the MixUp and TransMix baselines, MergeMix achieves significant performance improvements over multiple shuffling baseline models. For CAFormer-S12, we train for 200 epochs on the CIFAR100 dataset with the same settings as DeiT-Tiny. Table A7 shows that MergeMix obtained a second-best result, since CAFormer only has with Attention module of 2 stages (8 layers).
@@ D.2 Relationship between Mixing Ratios and Rewards
Figure A2: LLM-as-the-Judge results for different mixing ratios. Higher mixing ratios consistently yield higher reward scores across strong MLLMs. MergeMix aligns closely with these models, showing that the mixing ratio provides a meaningful preference signal.
To understand how the proposed ranking loss on synthetic mixed pairs approximates human preference learning, we conduct a LLM-as-the-judge-eval evaluation in which strong MLLMs score the mixed pairs generated with different mixing ratios $\lambda$. Specifically, we query a diverse set of frontier MLLMs, including Grok-3, Doubao-Seed-1.6 (guo2025seed1), Doubao-Seed-1.6-CoT, Qwen3-VL-Plus-32B (yang2025qwen3), Qwen2.5-VL-72B (bai2025qwen2.5), Gemini-2.5 (team2023gemini), and a human-expert baseline, to assign reward scores to each mixed pair. As shown in Figure A2, we found that the reward consistently increases with larger mixing ratios, and this trend holds across nearly all evaluators. The strong monotonic correlation suggests that $\lambda$ provides a reliable and well-behaved control signal for modeling preference strength. This observation validates our use of $\lambda$ as an interpretable proxy for preference supervision and highlights its potential as a lightweight alternative to explicit human-annotated reward signals.
@@ D.4 Results of Different Ranking Loss
mDOP (wang2024mdpo) 67.53 57.90 50.04 64.60 60.02 +0.20
Re-Align (xing2025re) 68.10 58.55 50.06 64.69 60.35 +0.53
@@ Appendix F Visualization and Case Study
In this section, we provide a visualization of the case study of augmentation samples with corresponding performance reward scores and an extensive visualization of mixing samples and ToMe source maps. Firstly, we provide a case of different degrees of augmentation in Figure A4. Then, we plot some visualizations of token merge with different merge ratios, mixed samples with different $\lambda$ in Figure A5 and Figure A6. For every three rows in Figure A5 and Figure A6, the first and second rows of source maps could directly capture the important regions of the raw images, where a large merge ratio enables better grouping of the similar regions. Based on the source maps, when the mixing ratio $\lambda$ grows from small to large, MergeMix could keep the most distinguishable tokens of the first image while gradually expanding more tokens from less important regions, which enables MergeMix to generate reliable mixing samples with no more computational costs. Moreover, we provide GradCAM (selvaraju2017gradcam) visualization of the top-1 and top-2 classes with the mixed samples of MergeMix on ImageNet-1K, as shown in Figure A7 and Figure A6.
@@
Figure A7: Visualization of mixed samples and corresponding GradCAM (selvaraju2017gradcam) of the top-1/2 class with MergeMix on ImageNet-1K.
Figure A8: Visualization of mixed samples and corresponding GradCAM (selvaraju2017gradcam) of the top-1/2 class with MergeMix on ImageNet-1K.

### v3

@@ 5.3 Implementations
In this subsection, we briefly introduce implementations on classification tasks and MLLM benchmarks (full details in Appendix B.2). For CIFAR100, ViT-based models (e.g., DeiT) use 224×224 images, trained with AdamW (weight decay 0.05), batch size 100, 200/600 epochs, RandomFlip/RandomCrop/RandAugment (Cubuk et al., 2020), and learning rates: 1e-3 (DeiT-Tiny/Small, cosine), 5e-4 (ViT-Small/Base), 2e-4 (ViT-Large). For ImageNet-1K, settings are consistent except 1e-3 learning rate, batch size 1024, 300 epochs (DeiT-Tiny/Small). For fine-grained datasets (CUB200, FGVC-Aircrafts, Stanford-Cars), DeiT-Small/ViT-Base are fine-tuned 200 epochs (batch size 16, 1e-5 lr) with PyTorch pretrained weights (Paszke et al., 2019).
Following LLaVA-v1.5, we use Vicuna-v1.5 7B (Chiang et al., 2023) (language decoder) and a 2-layer pre-trained MLP (visual-textual alignment, 1 epoch on LCS-558K), with a pre-trained CLIP vision encoder. SFT uses 1 epoch on llava-v1.5-mix665k (2e-5 lr, batch size 64, AdamW, 0.03 warmup, cosine scheduler), unfreezing the vision encoder (unlike LLaVA). For Qwen2.5-VL-Instruction, official checkpoints are fine-tuned 0.1 epoch on llava-v1.5-mix665k (similar settings); learning rates: 2e-6 (vision encoder), 1e-5 (LLM decoder/merger).
@@ 5.6 Results of Calibration
DNNs are prone to overconfidence in classification tasks. Mainfold Mixup (Verma et al., 2019) found that the mixup methods can effectively alleviate this problem. To this end, we compute the Expected Calibration Error (ECE (Guo et al., 2017)) of various mixup approaches on the CIFAR100 dataset for image classification. Also, to further analyze the calibration of MLLMs, we implement four short answer tasks, POPE, GQA, ScienceVQA, and SEED. Figure 4 shows the results of DeiT-Tiny and ViT-Small models trained for 200 epochs, showing that MergeMix obtains the best calibration of 6.7% and 9.7% in those ViT-specific mixup methods. (i.e., TransMix, SMMix, and MixPro). Table 6 shows the results of the LLaVA baseline, LLaVA with SFT, and LLaVA with our approach. SFT reduces the ECE when tuning the vision encoder, with augmentation and ranking loss, which can be better since we bring in the reward signal for the model. The more comprehensive results of CIFAR100 and MLLM benchmark we plot in Table A3 and Table A9.
@@ 6 Conclusion
Acknowledgement
This paper is supported by Young Scientists Fund of the National Natural Science Foundation of China (NSFC) (No. 62506305), Zhejiang Leading Innovative and Entrepreneur Team Introduction Program (No. 2024R01007), Key Research and Development Program of Zhejiang Province (No. 2025C01026), Scientific Research Project of Westlake University (No. WU2025WF003), Chinese Association for Artificial Intelligence (CAAI) & Ant Group Research Fund - AGI Track (No. 2025CAAI-ANT-13). It is also supported by the research funds of the National Talent Program and Hangzhou Municipal Talent Program.
@@ B.1 Datasets
We choose five mainstream classification datasets: (i) CIFAR100 dataset (Krizhevsky et al., 2009) consists of 100 classes of color images with a resolution of 32 $\times$ 32 pixels, containing 50,000 training images and 10,000 test images. (ii) ImageNet-1k dataset (Russakovsky et al., 2015) consists of 1,000 classes with varied image resolutions commonly cropped to 224 $\times$ 224 pixels, containing 1,281,167 training images and 50,000 test images. (iii) CUB200 dataset (Wah et al., 2011) contains 200 bird species, including total 11,788 images, we divided 5,994 images as training images and 5,794 images as test images, (iv) FGVC-Aircrafts dataset (Maji et al., 2013) contains 100 aircraft model classes, including 6,667 training images and 3,333 test images, (v) Stanford-Cars dataset (Krause et al., 2013) contains 196 car model classes, including 8,144 training images and 8,041 test images. All the fine-grained datasets we used set the resolution as 224 $\times$ 224 pixels for training and testing.
@@ B.2 Implementations
(i) For the CIFAR100 dataset, aiming to be suitable for training ViT-based approaches, e.g., DeiT, we resize images to $224\times 224$ and train them with the AdamW optimizer with weight decay of 0.05, batch size of 100, and total training of 200 epochs and 600 epochs. Uses RandomFlip and RandomCrop as basic augmentations, and additionally, we use RandAugment (Cubuk et al., 2020). For DeiT-Tiny and DeiT-Small, we use the learning rate of 1e-3 with a dynamic cosine scheduler. For ViT-Small and ViT-Base models, we set the learning rate to 5e-4, the learning rate of ViT-Large up to 2e-4, all dynamically adjusted by a cosine scheduler. (ii) For the ImageNet-1K dataset, the dataset settings are the same as CIFAR100, but we use the learning rate as 1e-3, batch size of 1024, and a total training of 300 epochs for DeiT-Tiny and DeiT-Small with AdamW optimizer with weight decay of 0.05. (iii) For all fine-grain datasets, i.e., CUB-200 dataset, FGVC-Aircrafts dataset, and Stanford-Cars dataset, we fine-tune the DeiT-Small and ViT-Base model for 200 epochs with a batch size of 16, learning rate of 1e-5, loading the pre-trained model weight from PyTorch (Paszke et al., 2019).
Table A1: Top-1 accuracy (%) of mixup methods on CIFAR-100 dataset under DeiT-Tiny/Small, ViT-Small/Base/Large different model sizes. The $\alpha$ parameter of the Beta distribution follows the setting in OpenMixup (Li et al., 2022) setting.
@@
Following the LLaVA-v1.5 settings, we use a pre-trained Vicuna-v1.5 7B (Chiang et al., 2023) as the language decoder, which uses a pre-trained 2 $\times$ MLP as the projection for aligning the vision and text modistes, which was trained for one epoch on LCS-558K. For the vision encoder, we use a pre-trained CLIP encoder and extract the visual representation from the input images. For SFT, the learning rate was set as 2e-5, the batch size was 64, and training one epoch on llava-v1.5-mix665k dataset, uses AdamW optimizer with (0.9, 0.999) betas and epsilon of 1e-8, warmup ratio of 0.03 with a cosine scheduler. The difference from LLaVA is that we unfreeze the vision encoder during training. About Qwen2.5-VL-Instruction, we fine-tune with llava-v1.5-mix665k dataset for 0.1 epoch, uses AdamW optimizer with (0.9, 0.999) betas and epsilon of 1e-8 like LLaVA, warmup with 0.03 ratio. The learning rate of the vision encoder, LLM decoder, and merger were set to 2e-6, 1e-5, and 1e-5, respectively.
@@ B.4 Future work
There are still some shortcomings in MergeMix for MLLMs. In future work, we will explore this from: (1) From the data perspective: MergeMix focuses on the enhancement of the image modality during training, while text inputs still remain raw. How to extend mixup to the text modality in MLLM tasks needs to be solved, as this can provide more fine-grained optimization guidelines. (2) From the model perspective: The token merging is still static and unlearned. Improving the merging strategy to make it learnable via metrics or backpropagation could enhance the token merging ability for mixing. (3) How to use sampling to find the optimal candidate as the loser from multiple augmented data, like RL-based methods (Feng et al., 2025b; a) and how to construct high-quality losers when adopting image editing methods (Wang et al., 2026) are also promising ways to obtain the preference pairs.
@@ C.2 Robustness Experiments of Mixup Augmentations
For further exploring the effectiveness of MergeMix, we applied our approach on a hyper-model MetaFormer (CAFormer) (Yu et al., 2022) and two audio datasets for classification by HuBERT-Base (Hsu et al., 2021). ESC-50 (Piczak, 2015) consisted of 50 classes, with 1,200 training samples and 400 validation samples, and a maximum duration of 3 seconds. UrbanSound8k (Salamon et al., 2014) is a classification dataset consisting of 10 classes, with a maximum duration of 4 seconds, containing 7,079 training samples and 816 validation samples. Following the USB experimental setup, we fine-tuned the model for 100 epochs using the AdamW optimizer. The base learning rate was set to 1e-4, texttt5e-4, with a batch size of 32 and a weight decay coefficient of 5e-4. Both the compared shuffling methods and our proposed MergeMix can be directly transferred to audio data (treated as one-dimensional sequences). Table A8 below shows our reproduced comparison results on two audio datasets. Compared to the MixUp and TransMix baselines, MergeMix achieves significant performance improvements over multiple shuffling baseline models. For CAFormer-S12, we train for 200 epochs on the CIFAR100 dataset with the same settings as DeiT-Tiny. Table A7 shows that MergeMix obtained a second-best result, since CAFormer only has with Attention module of 2 stages (8 layers).
@@ D.2 Relationship between Mixing Ratios and Rewards
Figure A2: VLM-as-the-Judge results for different mixing ratios. Higher mixing ratios consistently yield higher reward scores across strong MLLMs. MergeMix aligns closely with these models, showing that the mixing ratio provides a meaningful preference signal.
To understand how the proposed ranking loss on synthetic mixed pairs approximates human preference learning, we conduct a VLM-as-the-judge-eval evaluation in which strong MLLMs score the mixed pairs generated with different mixing ratios $\lambda$. Specifically, we query a diverse set of frontier MLLMs, including Grok-3, Doubao-Seed-1.6 (Guo et al., 2025), Doubao-Seed-1.6-CoT, Qwen3-VL-Plus-32B (Yang et al., 2025a), Qwen2.5-VL-72B (Bai et al., 2025), Gemini-2.5 (Team et al., 2023), and a human-expert baseline, to assign reward scores to each mixed pair. As shown in Figure A2, we found that the reward consistently increases with larger mixing ratios, and this trend holds across nearly all evaluators. The strong monotonic correlation suggests that $\lambda$ provides a reliable and well-behaved control signal for modeling preference strength. This observation validates our use of $\lambda$ as an interpretable proxy for preference supervision and highlights its potential as a lightweight alternative to explicit human-annotated reward signals.
@@ D.4 Results of Different Ranking Loss
mDOP (Wang et al., 2024a) 67.53 57.90 50.04 64.60 60.02 +0.20
Re-Align (Xing et al., 2025) 68.10 58.55 50.06 64.69 60.35 +0.53
@@ Appendix F Visualization and Case Study
In this section, we provide a visualization of the case study of augmentation samples with corresponding performance reward scores and an extensive visualization of mixing samples and ToMe source maps. Firstly, we provide a case of different degrees of augmentation in Figure A4. Then, we plot some visualizations of token merge with different merge ratios, mixed samples with different $\lambda$ in Figure A5 and Figure A6. For every three rows in Figure A5 and Figure A6, the first and second rows of source maps could directly capture the important regions of the raw images, where a large merge ratio enables better grouping of the similar regions. Based on the source maps, when the mixing ratio $\lambda$ grows from small to large, MergeMix could keep the most distinguishable tokens of the first image while gradually expanding more tokens from less important regions, which enables MergeMix to generate reliable mixing samples with no more computational costs. Moreover, we provide GradCAM (Selvaraju et al., 2019) visualization of the top-1 and top-2 classes with the mixed samples of MergeMix on ImageNet-1K, as shown in Figure A7 and Figure A8.
@@
Figure A7: Visualization of mixed samples and corresponding GradCAM (Selvaraju et al., 2019) of the top-1/2 class with MergeMix on ImageNet-1K.
Figure A8: Visualization of mixed samples and corresponding GradCAM (Selvaraju et al., 2019) of the top-1/2 class with MergeMix on ImageNet-1K.